PULSED FIELD IONIZATION ELECTRON SPECTROSCOPY OF CYCLOPENTADIENYL DIALUMINUM COMPLEX

\underline{YUXIU LEI} and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The cyclopentadienyl dialuminum complex, (C_{5}H_{5})Al_{2}, is synthesized in a supersonic molecular beam, where aluminum atoms and dimers are produced by laser vaporization, and cyclopentadienyl radicals are generated from photolysis of cyclopentadiene (C_{5}H_{6}) molecules. Vibrationally resolved electronic spectra of the complex are measured using pulsedfield ionization zero electron kinetic energy spectroscopy. Three initial conformations are considered for the complex: one with an Al atom binding to each side of the C_{5}H_{5} ring and two others with an Al_{2} molecule lying vertically or horizontally on one side of the ligand ring. A structure with the Al_{2} molecule lying vertically on the top of the C_{5}H_{5} ring is identified by comparing the spectroscopic measurements to ab initio calculations. The (C_{5}H_{5})Al_{2} neutral complex has a ^{2}A′′ ground electronic state in Jahn-Teller distorted C_{s} symmetry and the corresponding ion has a ^{1}A_{1} ground state in C_{5v} symmetry. The ionization energy of (C_{5}H_{5})Al_{2} is measured to be 40690(5) cm^{−1}. For the cationic ^{1}A_{1} state, the frequencies of the metal-ligand stretching, metal-metal stretching, and metal-ligand bending modes are 482, 175, and 50 cm^{−1}, respectively. In addition, a (C_{5}H_{5})-Al_{2} bending mode of the neutral ^{2}A′′ state is measured to be 26 cm^{−1}.